### 09 Mar 2026 v1

@@ 4.1 The Main Result
Proof. First, let us establish that the PSPE with $m^{\ast}=M$ is dynamically stable. By Lemma 1, for every $m\geq M,$ if there exists a solution $x_{m}>0$ to the equation $\phi_{m}(c_{m}x_{m})=x_{m},$ then it is unique. Moreover, $x_{m}^{\prime}>\phi_{m}(c_{m}x_{m}^{\prime})$ for every $x_{m}^{\prime}>x_{m}$ while $x_{m}^{\prime}<\phi_{m}(c_{m}x_{m}^{\prime})$ for every $x_{m}^{\prime}<x_{m}.$ Hence, for any such solution $x_{m}$,
$\phi_{m}(c_{m}x_{m}^{\prime})-\phi_{m}(c_{m}x_{m})=\phi_{m}(c_{m}x_{m}^{\prime})-x_{m}<x_{m}^{\prime}-x_{m}$
for any $x_{m}^{\prime}>x_{m}$, while
$\phi_{m}(c_{m}x_{m}^{\prime})-\phi_{m}(c_{m}x_{m})=\phi_{m}(c_{m}x_{m}^{\prime})-x_{m}>x_{m}^{\prime}-x_{m}$
for any $x_{m}^{\prime}<x_{m}$. Now consider $m<M$. By the definition of $M$ (as the lowest value of $m^{\prime}$ for which $\theta_{m^{\prime}}>0$), the only non-negative solution to $\phi_{m}(c_{m}x_{m})=x_{m}$ is $x_{m}=\phi_{m}(c_{m}x_{m})=0$; and by Lemma 1, $\phi_{m}(c_{m}x_{m}^{\prime})-x_{m}^{\prime}<0$ for all $x_{m}^{\prime}>0$. Therefore,
$\phi_{m}(c_{m}x_{m}^{\prime})-\phi_{m}(c_{m}x_{m})=\phi_{m}(c_{m}x_{m}^{\prime})<x_{m}^{\prime}=x_{m}^{\prime}-x_{m}$
@@
Finally, consider a PSPE with $m^{\ast}>M$. Thus, $x_{M}=0$. By the definition of $M$ and Lemma 1, we can find $x_{m}^{\prime}>0$ such that $\phi_{m}(c_{m}x_{m}^{\prime})-x_{m}^{\prime}>0$, which means that the equilibrium is not dynamically stable. $\blacksquare$
Thus, dynamic stability uniquely selects the PSPE with the lowest amount of offensive speech.
@@
From a narrowly technical point of view, what makes the equilibrium analysis of whataboutism distinct from the benchmark model, is that the cost of playing $a=1$ is non-linear in the behavior of others. This non-linearity arises from the way the success rate of a whatboutism rebuttal depends on the equilibrium state-dependent frequency of internal condemnation.
@@ 4.2 The Frequency of Offensive Speech and Whataboutism
The interpretation of the condition $\theta_{m}=\lambda b_{m}-2g_{m}+2>0$ (which holds for $m\geq M$) is that the probability of external condemnation ($\lambda b_{m}$) is more than twice the fraction of “fanatics” in the camp choosing whether to engage in offensive speech ($g_{m}-1$). The following result applies comparative statics to the expression (4.2).
@@
where the left-hand equality is obtained by plugging (4.2). This implies the following observation.
@@
Differentiate the equilibrium expressions for $x_{m}$ and the frequency of whatboutism in state $m$ w.r.t $k$, and evaluate the derivative at $k=1$. The derivative of $x_{m}$ is negative while the derivative of $(1-\theta_{m})^{2}$ is positive. This means that as society becomes more polarized, both offensive speech and whatboutism become more frequent.

### 16 Mar 2026 v2

@@ 4.1 The Main Result
Proof. We begin by establishing relations that are related to those given by Lemma 1. Suppose $\theta_{m}>0$. Denoting $w=c_{m}z$ and applying Lemma 1, we obtain that for any $w>0$: $c_{m}\phi_{m}(w)>w$ if $w<c_{m}\theta_{m}$; $c_{m}\phi_{m}(w)<w$ if $w>c_{m}\theta_{m}$; and $c_{m}\phi_{m}(w)=w$ has the unique solution $w=c_{m}\theta_{m}$. Now suppose $\theta_{m}\leq 0.$ Then, by the same argument, the only non-negative solution to $c_{m}\phi_{m}(w)=w$ is $w=0$, and $c_{m}\phi_{m}(w)<w$ for any $w>0$.
We now show that the PSPE with $m^{\ast}=M$ is dynamically stable. By the previous paragraph, for every $m\geq M,$ if there exists a solution $x_{m}>0$ to the equation $c_{m}\phi_{m}(x_{m})=x_{m},$ then it is unique. Moreover, $x>c_{m}\phi_{m}(x)$ for every $x>x_{m}$ while $x<c_{m}\phi_{m}(x)$ for every $x<x_{m}.$ Hence, for any such solution $x_{m}$,
$c_{m}\phi_{m}(x)-c_{m}\phi_{m}(x_{m})=c_{m}\phi_{m}(x)-x_{m}<x-x_{m}$
for any $x>x_{m}$, while
$c_{m}\phi_{m}(c_{m}x)-c_{m}\phi_{m}(x_{m})=c_{m}\phi_{m}(c_{m}x)-x_{m}>x-x_{m}$
for any $x<x_{m}$.
Next consider $m<M$. By the definition of $M$ (as the lowest value of $m^{\prime}$ for which $\theta_{m^{\prime}}>0$), and by the first paragraph, the only non-negative solution to $c_{m}\phi_{m}(x_{m})=x_{m}$ is $x_{m}=0$, and for any $x>0,$
$c_{m}\phi_{m}(x)-c_{m}\phi_{m}(x_{m})=c_{m}\phi_{m}(x)<x=x-x_{m}$
@@
Finally, consider a PSPE with $m^{\ast}>M$. By the definition of $m^{\ast}$, $x_{M}=0$. By the definition of $M,$ $\theta_{M}>0.$ Hence, by the first paragraph, for any $x\in(0,c_{M}\theta_{M}),$
$c_{M}\phi_{M}(x)-c_{M}\phi_{M}(x_{M})=c_{M}\phi_{M}(x)>x=x-x_{M}$
implying that the equilibrium is not dynamically stable. $\blacksquare$
It follows that dynamic stability uniquely selects the PSPE with the lowest amount of offensive speech. In the two-state example we analyzed earlier in this section, this has the following simple implication: If an equilibrium with $v^{\ast}>0$ exists, it is the unique stable one. Otherwise, the unique equilibrium $v^{\ast}=0$ is also stable.
@@
From a narrowly technical point of view, what makes the equilibrium analysis of whataboutism distinct from the benchmark model, is that the cost of playing $a=1$ is non-linear in the behavior of others. This non-linearity arises from the way the success rate of a whataboutism rebuttal depends on the equilibrium state-dependent frequency of internal condemnation.
@@ 4.2 The Frequency of Offensive Speech and Whataboutism
The interpretation of the condition $\theta_{m}=\lambda b_{m}-2g_{m}+2>0$ (which holds for $m\geq M$) is that the probability of external condemnation ($\lambda b_{m}$) is more than twice the fraction of “fanatics” in the camp choosing whether to engage in offensive speech ($g_{m}-1$). The following result applies comparative statics to the expression (13).
@@
where the left-hand equality is obtained by plugging (13). This implies the following observation.
@@
Differentiate the equilibrium expressions for $x_{m}$ and the frequency of whataboutism in state $m$ w.r.t $k$, and evaluate the derivative at $k=1$. The derivative of $x_{m}$ is negative while the derivative of $(1-\theta_{m})^{2}$ is positive. This means that as society becomes more polarized, both offensive speech and whataboutism become more frequent.